If the height of a triangle is twice its base and area of the triangle is 225cm2 , then what are its base and   height    .


1
Expert's answer
2021-03-11T12:30:16-0500

"S = \frac{1}{2} \cdot h \cdot a" , where "S" - area of triangle, "h" - height of triangle, "a" - base of triangle;

"S = 225" (cm2); "h = 2a" by condition; "\Rightarrow" After all replacements: "225 = \frac{1}{2} \cdot 2a \cdot a;"

"a^2 = 225" ; "a = 15" (cm) - value of base of triangle;

"h = 2a =2 \cdot 15 = 30" (cm) - value of height of triangle;

Answer : "a = 15" cm, "h = 30" cm.